# Symmetry in Neural Network Parameter Spaces

### v1

@@ 1 Introduction
Despite their remarkable empirical success, neural networks remain not fully understood from a theoretical standpoint. Designing effective architectures often relies on heuristics, and training outcomes can be highly sensitive to initialization and optimization details. The rapid growth of model sizes—exemplified by the recent explosion in the size of language and other generative models (Kaplan et al., 2020)—further complicates efforts to characterize training dynamics and generalization behavior. While theoretical work has clarified aspects of overparameterization (Neyshabur et al., 2018) and implicit regularization in gradient descent (Soudry et al., 2018), we still lack a unified understanding of how the structure of parameter space shapes the loss landscape and affects optimization.
@@ 2 Parameter Space Symmetry
Symmetry is prevalent in deep learning architectures. In this section, we explore various definitions and examples of parameter space symmetry, starting with transformations of the parameters that preserve the feedforward function. Then we expand our discussion to include relaxed definitions of symmetry, which preserve the overall loss function or the feedforward function’s value on subsets of data. Along with these definitions, we provide a list of known symmetries and show how they arise from the architecture of neural networks. Finally, we discuss the impact of symmetries on parameter identifiability and whether the known symmetries are complete.
@@ 2.2.1 Examples: Symmetries in Common Neural Network Components
In a radial basis function (RBF) network $\sum_{i=1}^{k}w_{i}\varphi\left(\frac{||\mathbf{x}-\mathbf{c_{i}}||}{b_{i}}\right)$ with radial function $\varphi\colon{\mathbb{R}}_{+}\to{\mathbb{R}}$ and parameters $(w_{i},b_{i},c_{i})_{i=1}^{k}\in\Theta=({\mathbb{R}}\times{\mathbb{R}}_{+}\times{\mathbb{R}}^{n})^{k}$, there is a permutation symmetry acting on $\Theta$ by $\pi\cdot((w_{i},b_{i},c_{i})_{i=1}^{k})=(w_{\pi^{-1}(i)},b_{\pi^{-1}(i)},c_{\pi^{-1}(i)})_{i=1}^{k}$, for $\pi\in\text{S}_{k}$ (Kůrková and Neruda, 1994).
@@
Radial basis function  $\sum_{i=1}^{k}w_{i}\varphi\left(\frac{||\mathbf{x}-\mathbf{c_{i}}||}{b_{i}}\right)$ $S_{k}$ $\pi\cdot(w_{i},b_{i},c_{i})=(w_{\pi^{-1}(i)},b_{\pi^{-1}(i)},c_{\pi^{-1}(i)})$
@@ 2.2.2 Example: Transformers
In addition to the $(\operatorname{GL}_{d_{k}}({\mathbb{R}}))^{h}$ symmetry inherited from individual attention heads, a multi-head attention admits an $S_{h}$ and a $(\operatorname{GL}_{d_{v}}({\mathbb{R}}))^{h}$ symmetry. The $S_{h}$ symmetry acts as permutation of attention heads: $\pi\cdot(W_{i}^{Q},W_{i}^{K},W_{i}^{V},W_{i}^{O})=(W_{\pi^{-1}(i)}^{Q},W_{\pi^{-1}(i)}^{K},W_{\pi^{-1}(i)}^{V},W^{O}_{\pi^{-1}(i)})$, for $\pi\in S_{h}$. The multiplication of each attention head with $W_{O}$ results in one $\operatorname{GL}_{d_{v}}({\mathbb{R}})$ symmetry, acting by $g_{i}\cdot(W_{i}^{Q},W_{i}^{K},W_{i}^{V},W_{i}^{O})=(W_{i}^{Q},W_{i}^{K},W_{i}^{V}g_{i}^{-1},g_{i}W^{O}_{i})$, for $g_{i}\in\operatorname{GL}_{d_{v}}({\mathbb{R}})$, $i=1,...,h$.
@@ 2.3.3 Distribution Symmetry
Consider the function $F\colon{\mathbb{R}}^{m\times n}\times({\mathbb{R}}^{n}\times{\mathbb{R}}^{n})\to{\mathbb{R}}^{m}$ defined by $F(W,(x,y))=Wx-y$, where $W$ is the parameter and $x,y$ are data. Assume that $x$ is drawn from a distribution $D_{x}$ and for each data pair $(x,y)$, $y$ is defined as $y=W^{*}x$ for a fixed matrix $W^{*}\in{\mathbb{R}}^{m\times n}$. Let $\bar{x}={\mathbb{E}}_{x\sim D_{x}}[x]$. Then, the expectation of $F$ under the data distribution is:
@@ 2.4 Parameter Identifiability and Completeness of Symmetry
The identifiability of ReLU networks is less straightforward. In particular, ReLU networks can contain hidden symmetries—transformations that preserve the realized function but not captured by permutation and positive scaling symmetries. Grigsby et al. (2023) describe several mechanisms through which hidden symmetries can arise. For example, for a single ReLU neuron with scalar input $x$ and realization map $\rho\colon(a,b)\mapsto(x\mapsto\text{ReLU}(ax+b))$, the fiber of the constant function 0, $\rho^{-1}(0)$, contains $(0,0)$ and $(0,-1)$, which are not related by rescaling and permutation (Grigsby et al., 2022). To formalize this variability, Grigsby et al. (2022) introduce the functional dimension, defined as the rank of the Jacobian of the realization map $\rho$, which reflects the local dimensionality of the set of functions realized near a given parameter. They show that the functional dimension may not be constant even within the same fiber. This suggest a complicated structure of symmetries in ReLU networks.
@@ 3 Role of Symmetry in Loss Landscapes
The loss landscape—the graph of the loss function over the parameter space—plays a central role in understanding the optimization behavior of neural networks. Parameter symmetries and the lack of parameter identifiability increase the size of loss level sets. To illustrate, consider a two–layer linear network $f_{linear}(W_{2},W_{1})=W_{2}W_{1}X$, where $(W_{2},W_{1})\in\Theta={\mathbb{R}}^{m\times h}\times{\mathbb{R}}^{h\times n}$ and $X\in{\mathbb{R}}^{n\times k}$ is fixed input data. If $(W_{2}^{*},W_{1}^{*})$ is a global minimum, so are all points $(W_{2}^{*}g^{-1},gW_{1}^{*})$ with $g\in\operatorname{GL}_{h}({\mathbb{R}})$. This set forms a positive-dimensional manifold of zero loss (Figure 6).
@@ 3.1 Continuous Symmetry and Mode Connectivity
Continuous parameter space symmetries enlarge individual minima into connected sets. Formally, a symmetry is continuous if the symmetry group $G$ is a Lie group and the group action is continuous. The orbit of a point $\theta\in\Theta$ is the set $O(\theta)=\{\theta^{\prime}\in\Theta\mid\theta^{\prime}=g\cdot\theta\text{ forsome }g\in G\}$. When $G$ is a Lie group, the orbits are often positive-dimensional, so continuous symmetries generically create a positive-dimensional manifold lying entirely in a single loss level set.
Empirical work first revealed that independently trained networks can be joined by low-loss curves in parameter space (Garipov et al., 2018; Draxler et al., 2018). This observation, often referred to as mode connectivity (Garipov et al., 2018), has been extended to show that points found by stochastic gradient descent (SGD) are connected by multi-dimensional volumes (Benton et al., 2021). This insight has applications in model ensembling (Garipov et al., 2018; Benton et al., 2021; Benzing et al., 2022) and model averaging (Izmailov et al., 2018; Wortsman et al., 2022). Continuous symmetries provide a theoretical explanation for this phenomenon. The loss remains constant along each orbit, so minima discovered by SGD will automatically be connected whenever the group action is continuous and they can be mapped to each other by an element in the identity component of the group.
Parameter space symmetry imposes a structure on the loss level sets. The minimum of a neural network, which is the optimization target, is perhaps the most interesting level set. Investigating the connection between symmetry groups and the topology of minima may lead to better understanding of the loss landscape and explain the source of mode connectivity. When a level set is homeomorphic to the symmetry group through a group action, e.g. in linear networks with invertible weight matrices, the minima has the same topological properties, such as connectedness, as the symmetry group (Zhao et al., 2025). Zero-loss curves generated by continuous actions (Zhao et al., 2024b) and dimension bounds derived from orbit dimensions (Zhao et al., 2023) further illustrate how symmetry dictates flat directions. (Lengyel et al., 2020) visualizes the set of functionally equivalent parameters. In ReLU networks, the visualization demonstrates the expected permutation and scaling symmetry, as well as other structures which indicate more complex symmetries.
@@ 3.2 Discrete Symmetry and Structure of Minimum
Discrete symmetries do not connect minima continuously; instead they replicate them, creating many functionally identical copies throughout the parameter space. Permutation is the best-studied discrete symmetry. Brea et al. (2019) show that all permutations of a given hidden layer reside in the same loss level set, while Şimşek et al. (2021) prove that adding just one extra neuron per layer in a minimal tanh network merges these replicas into a single connected manifold. Extending this analysis, Farrugia-Roberts (2023) characterize functional equivalence classes for hyperbolic-tangent networks, and Pittorino et al. (2022) examine minima in the function space after quotienting out both permutation and scaling symmetries. Counting arguments reveal the combinatorial explosion such symmetries induce. Michelucci (2022) estimate that the number of permutation-equivalent minima grows factorially with width, underscoring how discrete symmetry complicates landscape exploration.
@@ 3.3 Removing Symmetry: Model Compression and Reduced Search Space
One major motivation for removing symmetry is to reduce the dimensionality of the parameter space without altering the function space. This operation, often known as model compression, is of critical importance in lowering storage cost, improving inference speed, and efficient deployment of machine learning models. Every symmetry implies a set of directions in which the loss is constant, stretching level sets into high-dimensional manifolds. Quotienting out these directions yields a more compact version of the loss landscape, where each point corresponds to a unique function. In the context of model compression, this means we can represent networks more efficiently. For example, Ganev et al. (2022) compress radial neural networks by factoring out orthogonal symmetries, and Sourek et al. (2021) exploit structural symmetries in computational graphs to merge redundant nodes. These methods reduce storage costs and inference time, without affecting expressivity or accuracy.
[…]
@@
Finally, removing symmetry alters the curvature and critical point structure of the landscape, with direct implications for optimization. Symmetry-induced flat directions often give rise to plateaus or saddle points that slow down training. By projecting the loss landscape onto a symmetry-reduced space, these degenerate directions are removed, yielding a more convex surface. Leake and Vishnoi (2021) show that continuous symmetries can be used to construct convex polytopes over which equivalent optimization problems become easier to solve.
@@ 4.1 Exploiting Difference among Points in an Orbit
Figure 7: Implications and applications of parameter space symmetry in gradient-based optimization. (a) Points in an orbit have same loss values $L$ but different gradients $\nabla L$. (b) Gradient descent trajectories originating from two points in the same orbit, $\theta$ and $g\cdot\theta$, may differ. (c) Algorithms leveraging these differences search within the orbit for a point that leads to faster optimization ($g^{*}\cdot\theta$). (d) A $G$-invariant algorithm $A$, which commutes with the group action, is indifferent to initializations within an orbit.
@@
Even when the group actions do not preserve loss exactly, the relaxed symmetry is a useful tool to explore the orbits. In this case, one can leverage the different loss or functions realized by points in the same orbit. In certain methods with regularizers, loss is not invariant on the orbits, but the variation is small, which slows down optimization after parameters enter the orbit. This is analogous to the Goldstone mode in physics (Altland and Simons, 2010). After optimization on the full parameter space slows down due to entering the Goldstone mode, Bamler and Mandt (2018) propose to optimize on the symmetry subspace, which accelerates optimization in these orbits with weakly-broken symmetry.
@@ 4.2.1 Scaling Invariant Algorithms
Some rescale invariant algorithms are based on the observation that the product of incoming and outgoing parameters of a ReLU function is invariant to rescaling. From the parameters $\theta$ of a ReLU network, Neyshabur et al. (2015) construct a path vector $\pi(\theta)$ that contains the product of parameters along each path from an input node to an output node. Their algorithm, Path-SGD, performs proximal gradient descent with respect to a path regularizer $\left\|\pi(\theta)-\pi(\theta^{t})\right\|_{p}^{2}$, instead of the 2-norm of parameters $\left\|\theta-\theta^{t}\right\|_{2}^{2}$ that would give rise to Equation 1:

### v2

@@ 1 Introduction
Despite their remarkable empirical success, neural networks remain not fully understood from a theoretical standpoint. Designing effective architectures often relies on heuristics, and training outcomes can be highly sensitive to initialization and optimization details. The rapid growth of model sizes—exemplified by the recent explosion in the size of language and other generative models (Kaplan et al., 2020)—further complicates efforts to characterize training dynamics and generalization behavior. While theoretical work has clarified aspects of overparameterization (Belkin et al., 2019; Neyshabur et al., 2018) and implicit regularization in gradient descent (Soudry et al., 2018), we still lack a unified understanding of how the structure of parameter space shapes the loss landscape and affects optimization.
@@ 2 Parameter Space Symmetry
In this section, we explore various definitions and examples of parameter space symmetry, starting with transformations of the parameters that preserve the feedforward function. Then we expand our discussion to include relaxed definitions of symmetry, which preserve the overall loss function or the feedforward function’s value on subsets of data. Along with these definitions, we provide a list of known symmetries and show how they arise from the architecture of neural networks. Finally, we discuss the impact of symmetries on parameter identifiability and whether the known symmetries are complete.
@@ 2.2.1 Examples: Symmetries in Common Neural Network Components
[…]
@@
[…]
@@ 2.2.2 Example: Symmetries in Transformers
Building on Example 2.10, multi-head attention replicates the attention block across $h$ heads and linearly mixes their outputs, thereby inheriting the $(\operatorname{GL}_{d_{k}}({\mathbb{R}}))^{h}$ symmetry and introducing additional head-permutation ($S_{h}$) and per-head linear symmetries ($(\operatorname{GL}_{d_{v}}({\mathbb{R}}))^{h}$).
@@
[…]
Parameters in a transformer can be finetuned after training to adapt to new datasets or tasks. A commonly used method, low-rank adaptation (LoRA) (Hu et al., 2022), updates pretrained parameters by low-rank matrices to reduce memory and compute. This parametrization adopts a general linear symmetry group.
Example 2.12 (LoRA).
In a low-rank adaptation $W+UV$, with pretrained weight matrix $W\in{\mathbb{R}}^{n\times m}$ and trainable parameters $U\in{\mathbb{R}}^{n\times r},V\in{\mathbb{R}}^{r\times m}$, there is a $\text{GL}_{r}({\mathbb{R}})$ symmetry, acting on $(U,V)$ by $g\cdot(U,V)=(Ug^{-1},gV)$, for $g\in\text{GL}_{r}({\mathbb{R}})$ (Putterman et al., 2024).
@@ 2.3.3 Distribution Symmetry
Consider the function $F\colon{\mathbb{R}}^{m\times n}\times({\mathbb{R}}^{n}\times{\mathbb{R}}^{n})\to{\mathbb{R}}^{m}$ defined by F(W,(x,y))=WxyF(W,(x,y))=Wx-y, where $W$ is the parameter and $x,y$ are data. Assume that $x$ is drawn from a distribution $D_{x}$ and for each data pair $(x,y)$, $y$ is defined as $y=W^{*}x$ for a fixed matrix $W^{*}\in{\mathbb{R}}^{m\times n}$. Let $\bar{x}={\mathbb{E}}_{x\sim D_{x}}[x]$. Then, the expectation of $F$ under the data distribution is:
@@ 2.4 Parameter Identifiability and Completeness of Symmetry
The identifiability of ReLU networks is less straightforward. In particular, ReLU networks can contain hidden symmetries—transformations that preserve the realized function but not captured by permutation and positive scaling symmetries. Grigsby et al. (2023) describe several mechanisms through which hidden symmetries can arise. For example, for a single ReLU neuron with scalar input $x$ and realization map $\rho\colon(a,b)\mapsto(x\mapsto\text{ReLU}(ax+b))$, the fiber of the constant function 0, $\rho^{-1}(0)$, contains $(0,0)$ and (0,1)(0,-1), which are not related by rescaling and permutation (Grigsby et al., 2022). To formalize this variability, Grigsby et al. (2022) introduce the functional dimension, defined as the rank of the Jacobian of the realization map $\rho$, which reflects the local dimensionality of the set of functions realized near a given parameter. They show that the functional dimension may not be constant even within the same fiber. This suggest a complicated structure of symmetries in ReLU networks.
@@
The identifiability of neural networks that are polynomial in their parameters can often be analyzed using tools from algebraic geometry (Marchetti et al., 2025). A notable example is multi-layer perceptrons with polynomial activation functions, or polynomial neural networks. Kileel et al. (2019) conjecture that permutation and scaling are the only symmetries in generic polynomial neural networks, which has been proven in the monomial case (Finkel et al., 2024). For more general polynomial activation functions, Shahverdi et al. (2025) show that there are only finitely many data-independent parameter symmetries. Moreover, they prove that convolutional neural networks with polynomial activations generically have no nontrivial data-independent parameter symmetries. Another example of a polynomial function is lightening self-attention $(W^{Q},W^{K},W^{V})\mapsto QW^{Q}(KW^{K})^{T}VW^{V}$, where the softmax normalization is omitted from the standard attention mechanism (Example 2.10). (Henry et al., 2025) prove that in a deep network composed of lightning self-attention layers, the only symmetries are (1) scaling $W_{Q}W_{K}^{T}$ and $W_{V}$ by a constant, (2) scaling $W_{Q}$ and $W_{K}$ by an invertible matrix, and (3) scaling $W_{V}$ of one layer and $W_{Q},W_{K}$ in the next layer by an invertible matrix. They conjecture and verify numerically that adding back the softmax breaks only the first of these symmetries.
@@ 3 Role of Symmetry in Loss Landscapes
The loss landscape—the graph of the loss function over the parameter space—plays a central role in understanding the optimization behavior of neural networks. Parameter symmetries and the lack of parameter identifiability increase the size of loss level sets. To illustrate, consider a two–layer linear network $f_{linear}(W_{2},W_{1})=W_{2}W_{1}X$, where $(W_{2},W_{1})\in\Theta={\mathbb{R}}^{m\times h}\times{\mathbb{R}}^{h\times n}$ and $X\in{\mathbb{R}}^{n\times k}$ is fixed input data. If $(W_{2}^{*},W_{1}^{*})$ is a global minimum, so are all points $(W_{2}^{*}g^{-1},gW_{1}^{*})$ with $g\in\operatorname{GL}_{h}({\mathbb{R}})$. This set forms a positive‐dimensional manifold of zero loss (Figure 6).
@@ 3.1 Continuous Symmetry and Mode Connectivity
Continuous parameter space symmetries enlarge individual minima into connected sets. Formally, a symmetry is continuous if the symmetry group $G$ is a Lie group and the group action is continuous. The orbit of a point $\theta\in\Theta$ is the set $O(\theta)=\{\theta^{\prime}\in\Theta\mid\theta^{\prime}=g\cdot\theta\text{ for some }g\in G\}$. When $G$ is a Lie group, the orbits are often positive-dimensional, so continuous symmetries generically create a positive‑dimensional manifold lying entirely in a single loss level set.
Empirical work first revealed that independently trained networks can be joined by low‑loss curves in parameter space (Garipov et al., 2018; Draxler et al., 2018). This observation, often referred to as mode connectivity (Garipov et al., 2018), has been extended to show that points found by stochastic gradient descent (SGD) are connected by multi-dimensional volumes (Benton et al., 2021). This insight has applications in model ensembling (Garipov et al., 2018; Benton et al., 2021; Benzing et al., 2022) and model averaging (Izmailov et al., 2018; Wortsman et al., 2022). Continuous symmetries provide a theoretical explanation for this phenomenon. The loss remains constant along each orbit, so minima discovered by SGD will automatically be connected whenever the group action is continuous and they can be mapped to each other by an element in the identity component of the group. We note that not all points in a minimum are related by a symmetry transformation, and empirical connectivity observed between independently trained networks may also arise from other mechanisms beyond symmetry.
Parameter space symmetry imposes a structure on the loss level sets. The minimum of a neural network, which is the optimization target, is perhaps the most interesting level set. Investigating the connection between symmetry groups and the topology of minima may lead to better understanding of the loss landscape and explain the source of mode connectivity. When a level set is homeomorphic to the symmetry group through a group action, e.g. in linear networks with invertible weight matrices, the minima has the same topological properties, such as connectedness, as the symmetry group (Zhao et al., 2025). Zero‑loss curves generated by continuous actions (Zhao et al., 2024b) and dimension bounds derived from orbit dimensions (Zhao et al., 2023) further illustrate how symmetry dictates flat directions. (Lengyel et al., 2020) visualizes the set of functionally equivalent parameters. In ReLU networks, the visualization demonstrates the expected permutation and scaling symmetry, as well as other structures which indicate more complex symmetries.
@@ 3.2 Discrete Symmetry and Structure of Minimum
Discrete symmetries do not connect minima continuously; instead they replicate them, creating many functionally identical copies throughout the parameter space. Permutation is the best‑studied discrete symmetry. Brea et al. (2019) show that all permutations of a given hidden layer reside in the same loss level set, while Şimşek et al. (2021) prove that adding just one extra neuron per layer in a minimal tanh network merges these replicas into a single connected manifold. Extending this analysis, Farrugia-Roberts (2023) characterize functional equivalence classes for hyperbolic‑tangent networks, and Pittorino et al. (2022) examine minima in the function space after quotienting out both permutation and scaling symmetries. Counting arguments reveal the combinatorial explosion such symmetries induce. Michelucci (2022) estimate that the number of permutation‑equivalent minima grows factorially with width, underscoring how discrete symmetry complicates landscape exploration.
@@ 3.3 Removing Symmetry: Model Compression and Reduced Search Space
One major motivation for removing symmetry is to reduce the dimensionality of the parameter space without altering the function space. This operation, often known as model compression, is of critical importance in lowering storage cost, improving inference speed, and efficient deployment of machine learning models. Although most production pipelines use pruning, quantization, or distillation, symmetry-aware reductions offer a function-preserving complement that can be combined with these methods. Every symmetry implies a set of directions in which the loss is constant, stretching level sets into high-dimensional manifolds. Quotienting out these directions yields a more compact version of the loss landscape, where each point corresponds to a unique function. In the context of model compression, this means we can represent networks more efficiently. For example, Ganev et al. (2022) compress radial neural networks by factoring out orthogonal symmetries, and Sourek et al. (2021) exploit structural symmetries in computational graphs to merge redundant nodes. These methods reduce storage costs and inference time, without affecting expressivity or accuracy.
Beyond compression, removing symmetry transforms the loss landscape by collapsing symmetry-related regions into single representatives. For networks with permutation symmetry—such as feedforward networks with interchangeable hidden units—the parameter space can be partitioned into large equivalence classes. Building on the observation that permutations of neurons can be viewed as compositions of reflections in the parameter space, Hecht-Nielsen (1990) constructs a cone in the parameter space that contains a permuted copy of every point in the parameter space. For two layer networks with permutation symmetry $S_{h}$, the cone occupies only $1/h$ of the parameter space. They further prove the existence of a minimal search set, in which no two points are related by symmetry. This geometric reduction reduces the volume of minima and reveals its global structure. Lim et al. (2024b) empirically show that after removing permutation and scaling symmetries, neural networks are more linearly mode connected, and the loss decreases more monotonically on the linear interpolation between initialization and trained parameters.
@@
Finally, removing symmetry alters the curvature and critical point structure of the landscape, with direct implications for optimization. Symmetry-induced flat directions often give rise to plateaus or saddle points that slow down training. By projecting the loss landscape onto a symmetry-reduced space, these degenerate directions are removed, yielding a more strongly convex surface. Leake and Vishnoi (2021) show that continuous symmetries can be used to construct convex polytopes, allowing nonconvex optimization problems to be reformulated as linear optimization problems over polytopes, which are computationally more tractable.
@@ 4.1 Exploiting Difference among Points in an Orbit
Figure 7: Implications and applications of parameter space symmetry in gradient-based optimization. (a) Points in an orbit have same loss values $L$ but different gradients $\nabla L$. (b) Gradient descent trajectories originating from two points in the same orbit, $\theta$ and $g\cdot\theta$, may differ. (c) Algorithms leveraging these differences search within the orbit for a point that leads to faster optimization ($g^{*}\cdot\theta$). (d) A $G$-invariant algorithm AA, which commutes with the group action, is indifferent to initializations within an orbit.
@@
Group actions that preserve loss approximately but not exactly is still a useful tool to explore near a loss level set. In this case, one can leverage the different loss or functions realized by points in the same orbit. In certain methods with regularizers, loss is not invariant on the orbits, but the variation is small, which slows down optimization after parameters enter the orbit. This is analogous to the Goldstone mode in physics (Altland and Simons, 2010). After optimization on the full parameter space slows down due to entering the Goldstone mode, Bamler and Mandt (2018) propose to optimize on the symmetry subspace, which accelerates optimization in these orbits with weakly-broken symmetry.
@@ 4.2.1 Scaling Invariant Algorithms
Two families of rescale invariant algorithms have been developed. The first one is based on the observation that the product of incoming and outgoing parameters of a ReLU function is invariant to rescaling. From the parameters $\theta$ of a ReLU network, Neyshabur et al. (2015) construct a path vector $\pi(\theta)$ that contains the product of parameters along each path from an input node to an output node. Their algorithm, Path-SGD, performs proximal gradient descent with respect to a path regularizer $\left\|\pi(\theta)-\pi(\theta^{t})\right\|_{p}^{2}$, instead of the 2-norm of parameters $\left\|\theta-\theta^{t}\right\|_{2}^{2}$ that would give rise to Equation 1: